### v1

# Negative Object Presence Evaluation (NOPE) to Measure Object Hallucination in Vision-Language Models

Abstract
Object hallucination poses a significant challenge in vision-language (VL) models, often leading to the generation of nonsensical or unfaithful responses with non-existent objects. However, the absence of a general measurement for evaluating object hallucination in VL models has hindered our understanding and ability to mitigate this issue. In this work, we present NOPE (Negative Object Presence Evaluation), a novel benchmark designed to assess object hallucination in VL models through visual question answering (VQA). We propose a cost-effective and scalable approach utilizing large language models to generate 29.5k synthetic negative pronoun ($\mathrm{NegP}$) data of high quality for NOPE. We extensively investigate the performance of 10 state-of-the-art VL models in discerning the non-existence of objects in visual questions, where the ground truth answers are denoted as $\mathrm{NegP}$(e.g., "none"). Additionally, we evaluate their standard performance on visual questions on 9 other VQA datasets. Through our experiments, we demonstrate that no VL model is immune to the vulnerability of object hallucination, as all models achieve accuracy below 10% on $\mathrm{NegP}$. Furthermore, we uncover that lexically diverse visual questions, question types with large scopes, and scene-relevant objects capitalize the risk of object hallucination in VL models.
1 Introduction
Figure 1: Example of object hallucination and incorrectness in VQA. The model hallucinates a non-existent man sitting on the closest bench in the left image, while in the right image, it simply answers inaccurately.
In recent years, vision-language (VL) research has witnessed a proliferation of studies focusing on diverse methods, models, and learning strategies aimed at bridging the performance gap between human and model capabilities Yang et al. (2021); Yi et al. (2018); Zhou et al. (2020); Ray et al. (2019); Gokhale et al. (2020); Dai et al. (2021, 2022); Ishii et al. (2021); Lovenia et al. (2022); Ji et al. (2022b); Lovenia et al. (2023). Furthermore, researchers have constructed more rigorous VL benchmarks to continually raise the performance standard Antol et al. (2015); Sheng et al. (2021); Li et al. (2021b); Goyal et al. (2017); Marino et al. (2019). However, despite these efforts, VL models continue to grapple with the persistent issue of object hallucination, where generated responses unfaithfully contain objects non-existent in the input images Ji et al. (2022a); Rohrbach et al. (2018); Dai et al. (2023b); Kayhan et al. (2021). As illustrated in Figure 1, the failure of the model to faithfully ground the visual input leads to the production of unfaithful answers. These instances of object hallucination not only result in incorrect responses but also shed light on fundamental issues within VL models, such as over-reliance on unimodal priors Jing et al. (2020); Agrawal et al. (2018); Gupta et al. (2022); Niu et al. (2021a) and statistical bias Agrawal et al. (2016); Goyal et al. (2017); Agarwal et al. (2020). These underlying problems impede the models’ ability to comprehend the concept of non-existence.
Despite the critical importance of addressing object hallucination in VL models, only a limited number of previous works have focused on mitigating this issue, primarily due to the challenges posed by the existing evaluation method in terms of generalization and scalability. CHAIR Rohrbach et al. (2018) has primarily concentrated on evaluating non-existent objects based on handcrafted parsing criteria as well as a predefined list of object categories and their synonyms in the context of image captioning tasks, typically utilizing 80 object categories from MSCOCO Rohrbach et al. (2018); Biten et al. (2022); Yi et al. (2018). However, the applicability of CHAIR to other datasets requires the generation of a new object category list, which exhibits varying levels of granularity across different studies Dai et al. (2023b); Biten et al. (2022).
In this paper, we present NOPE (Negative Object Presence Evaluation) to quantitatively assess object hallucination through VQA. We establish a clear distinction between object hallucination and incorrectness as follows: a) object hallucination refers to the phenomenon in VQA where a VL model’s response includes a non-existent object, despite the ground truth answer being a negative indefinite pronoun (e.g., "none", "no one", "nobody", "nowhere", "neither") Quirk et al. (1985) ($\mathbf{NegP}$); and b) incorrectness occurs when a VL model fails to accurately respond to a question with a ground truth answer that is anything other than $\mathrm{NegP}$, denoted as $\mathbf{Others}$ = $\mathbb{P}\setminus$$\mathrm{NegP}$, where $\mathbb{P}$ represents the set of all phrases. By leveraging $\mathrm{NegP}$, we evaluate object hallucination in NOPE, while $\mathrm{Others}$ allows us to assess normative correctness across diverse corpora. Our contributions are as follows:
  1. 1.
    By utilizing NOPE, we construct a VQA diagnostic benchmark to measure the object hallucination rate of VL models. Our experiment covers a balanced proportion of $\mathrm{NegP}$ and $\mathrm{Others}$ data with a total of $\sim$30k and $\sim$36k data in the dev and test sets, and includes 10 state-of-the-art VL baselines performances. We provide an in-depth analysis of the performances and limitations of the baselines.
  2. 2.
    We propose a novel automatic data generation pipeline to produce high-quality $\mathrm{NegP}$ VQA data from existing image captioning data by multi-turn prompting instruction-tuned large language models (LLMs). We verify and analyze our generated $\mathrm{NegP}$ data through automatic validation and human validation. Our list-then-rewrite method produces high-quality $\mathrm{NegP}$ VQA data with 92% validity.
  3. 3.
    Through extensive analysis in NOPE, we find that VL models tend to hallucinate more on data with higher lexical diversity, more scene-relevant objects, and larger answer scopes.
2 Related Work
2.1 Hallucination in Vision-Language
Only a few works study hallucination in vision-language, with the vast majority of them focusing on the task of image captioning.  Rohrbach et al. (2018) propose CHAIR, an automatic evaluation metric to measure object hallucination in generated image captions, which is defined as a phenomenon where the models produce captions containing objects that do not exist in the input visual context. Rohrbach et al. (2018); Dai et al. (2023b); Sharma et al. (2018) also show that standard captioning metrics, e.g., CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), SPICE Niu et al. (2022), under-penalize object hallucination. These evaluations open up a way for efforts to mitigate hallucination in image captioning Biten et al. (2022); Zhang et al. (2021); Xiao and Wang (2021); Dai et al. (2023b). Concurrent to our work, Li et al. (2023b) propose POPE and frame the task of evaluating object hallucination as a binary-class VQA with only "yes/no" answer.
2.2 Question Generation for VQA Data
Most works rely on human annotators to generate visual questions with ensured quality: VQAv2.0 and VQAv1.0 Goyal et al. (2017); Antol et al. (2015), Visual Genome Krishna et al. (2016), Visual7W Zhu et al. (2016), AdVQA Sheng et al. (2021), Vizwiz Gurari et al. (2018, 2019), TextVQA Singh et al. (2019), R-VQA Lu et al. (2018), VQA-Rephrasings Shah et al. (2019), etc.
However, the cost of human annotation is expensive, thus encouraging the exploration of a more scalable option: automatic VQA data generation.  Ren et al. (2015) present a simple question generation algorithm with a syntactic parser to convert image descriptions into QA forms. Johnson et al. (2017) use a functional program to generate synthetic images of objects as well as their relationships and relevant QA pairs using the ground-truth annotations. Kafle and Kanan (2017) populate multiple question templates with the image annotations (e.g., region descriptions, relationship graphs, bounding boxes) obtained from image captioning data to construct TDIUC. Changpinyo et al. (2022) annotate candidate answers by syntactically parsing the captions, then derive questions from them. While prior studies focus on generating $\mathrm{Others}$ VQA data, we aim to generate $\mathrm{NegP}$ VQA data, which has never been done by past works.
3 NOPE to Overcome Limited $\mathbf{NegP}$
Figure 2: Only 0.4% of existing VQA corpora consist of $\mathrm{NegP}$ data. The rest 99.6% is $\mathrm{Others}$.
As shown in Figure 2, there is only a minuscule amount of $\mathrm{NegP}$ data in the existing VQA datasets. In total, there are only $\sim$0.4of the existing VQA datasets are $\mathrm{NegP}$, which are not sufficient to assess object hallucination in VL. For this reason, we create NOPE through a novel $\mathrm{NegP}$ data generation method that aims to produce questions whose ground truth answers point to the absence of appropriate existent objects. Such ground truth $\mathrm{NegP}$ answers are denoted as $A^{\mathrm{NegP}}=\{"none","nothing","nowhere","zero","0",\\ "no\;one","nobody","neither"\}$. We automatically generate synthetic $\mathrm{NegP}$ VQA data by leveraging the zero-shot prompting abilities of pre-trained LLMs. To ensure the quality, we analyze the generated synthetic $\mathrm{NegP}$ VQA data through both automatic and manual human evaluation. The resulting $\mathrm{NegP}$ dataset is referred to as NOPE (Negative Object Presence Evaluation).
3.1 Prompting Methodology
We utilize an image captioning dataset $\mathcal{D}_{cap}=\{(v_{i},c_{i},l_{i}\}^{n}_{i=1}$, where $v_{i}$ denotes a visual context, $c_{i}$ denotes a textual caption, and $l_{i}$ denotes the relevant image label annotations (i.e., names of objects in $v_{i}$). We rely on $c_{i}$ to describe the objects and the relationship between objects depicted in $v_{i}$. We explore two prompting methods with varying degrees of flexibility to generate $\mathrm{NegP}$ questions from image captions: generate-from-scratch and list-then-rewrite. For clarity, we include all prompt templates with the examples in Appendix A and the automatic validation methods to ensure the validity of the generated questions in Appendix C.
Generate-from-scratch
In this method, we prompt an LLM to generate a question $q_{i}$ given three different variables: 1) an interrogative word $w_{i}\in\{"what","where","how\;many","who",\\ "which"\}$ to assert the question type needed for $q_{i}$, 2) a ground truth $\mathrm{NegP}$ answer $a_{i}\in A^{\mathrm{NegP}}$ that matches $w_{i}$, and 3) an image caption $c_{i}$.
Figure 3: Human evaluation results of $\mathrm{NegP}$ questions by generate-from-scratch and list-then-rewrite according to the categories in §3.2.
List-then-rewrite
LLMs can infer conversational contexts and follow instructions over multiple turns Nijkamp et al. (2023); Volum et al. (2022); Bang et al. (2023). Leveraging this multi-turn capability of LLMs, we frame our question generation task into two steps. (1) For object listing, given an image caption $c_{i}$ and the relevant object annotations $l_{i}$, we prompt an LLM to list $m$ objects $o_{i}=\{o_{i,j}\}^{m}_{j=1}$ that are “closely related” [^1] We use “closely related” (hard) for brevity. However, this object-scene relevance can be switched to “loosely related” or “completely unrelated” in practice. but not mentioned. (2) For question rewriting, the LLM has to paraphrase a provided reference question, which is sourced from a diverse pool of human-generated question templates with an object placeholder in Appendix B. After obtaining $m$ listed objects from (1), we pick $m$ random question templates from the pool and replace the object placeholders with the listed objects $o_{i}$ to construct the reference questions $r_{i}=\{r_{i,j}\}^{m}_{j=1}$. We prompt the LLM to paraphrase $r_{i}$ to $q_{i}=\{q_{i,j}\}^{m}_{j=1}$ to increase the lexical variety of the rewritten questions $q_{i}$.
3.2 Human Evaluation Guidelines
We conduct a human evaluation to verify and analyze the quality of the generated questions obtained from §3.1, as well as measure the effectiveness of the automatic validations performed. We employ three human annotators to perform the human evaluations. Detailed guidelines and examples are given prior to evaluation. We collect generated questions that are judged as valid and invalid by their automatic validation methods. Given a visual context, an image caption, a ground truth answer $\in A^{\mathrm{NegP}}$, and a generated question, the annotators are asked to determine whether: 1) the question is valid, 2) the question has a possible $\mathrm{Others}$ answer alternative, 3) the question does not match the answer (according to both the image caption and the image), 4) the question does not match the answer (only according to the image), or 5) the question is unclear or confusing. The examples provided for each category can be seen in Appendix E.
Figure 4: Distribution of NOPE’s $\mathrm{NegP}$ questions by their starting phrases. The arc length is proportional to the number of questions containing the word.
3.3 Results and Quality Analysis
Using automatic validation approaches explained in §3.1 and implementation details in Appendix D, we compare the capabilities of various instruction-tuned LLMs in generating $\mathrm{NegP}$ VQA data. From the automatic validation results and analysis presented in Appendix F, we find that employing ChatGPT yields the highest-quality generated $\mathrm{NegP}$ questions by both generate-from-scratch and list-then-rewrite prompting methods, hence its use in the human evaluations. We conduct a human evaluation on randomly selected 150 generated questions from each method. For each sample, we ask 3 human experts to judge each generated question into one of the 5 options defined in §3.2.
Figure 5: Object-scene relevance in the NOPE dataset. Related denotes “closely related” and unrelated denotes “completely unrelated” for brevity.
dev test
$\mathbf{NegP}$ 14718 17983
NOPE (§3.4) 14718 14773
AdVQA 0 88
R-VQA 0 9
TDIUC 0 6
Visual7W 0 1276
VQAv1 Abstract Scenes 0 180
VQAv2 Balanced Real 0 1651
dev test
$\mathbf{Others}$ 14850 18150
AdVQA 1350 1650
R-VQA 2700 3300
TDIUC 1350 1650
TextVQA 1350 1650
Visual7W 2700 3300
VizWiz 1350 1650
VQA-Rephrasings 1350 1650
VQAv1 Abstract Scenes 1350 1650
VQAv2 Balanced Real 1350 1650
Table 1: The data statistics of $\mathbf{NegP}$ (left) and $\mathbf{Others}$ (right) subsets used in the evaluation.
Figure 3 shows the result of our human evaluation. For generate-from-scratch, only $\pm$50out of the subset that is judged as valid by the automatic validation is actually a valid and appropriate $\mathrm{NegP}$ question according to the human annotators, and the rest is judged as incorrect by human annotators. The list-then-rewrite prompting approach, on the other hand, displays a significantly better question-answer generation quality with $\pm$92of the generated questions denoted as valid by the human annotators. This fact demonstrates that existing LLMs still fail to perform complex tasks in an end-to-end manner, while decomposing the complex tasks into several subtasks and coupling them with simple rule-based approaches can significantly improve the LLMs’ ability to perform the complex task effectively and efficiently.
A closer look at the questions generated by the generate-from-scratch method shows that while LLMs usually succeed in making questions in an end-to-end manner, 12% of the $\mathrm{NegP}$ generated questions include an existing object even though this information is sufficiently provided by the image caption. Moreover, 14% of the time, the generated questions also fail to include any objects and are overly generic, e.g., “What is not included in this image?”, which aligns with the observations of Jang et al. (2023); Hosseini et al. (2021); Ettinger (2020); Kassner and Schütze (2020) that LMs perform poorly on negation and struggles to understand that negation changes semantics. These facts show that LLMs cannot consistently perform this implicit task breakdown. From this human evaluation result, we can conjecture that the generate-from-scratch prompting method is not reliable and fails to elicit the LLMs’ understanding of complex tasks such as question generation. Using the list-then-rewrite method, we generate 29.5k $\mathrm{NegP}$ VQA data to build the NOPE dataset from OpenImagesV7 Kuznetsova et al. (2020).
3.4 Dataset Statistics
$\mathbf{NegP}$ Question Distribution
We cluster the generated questions into various types based on the starting n-grams in Figure 4. NOPE dataset exhibits a very broad lexical diversity of the generated questions, including variations in which the questions start with words other than the typical interrogative words (e.g., “what”, “where”, “how”, etc.), such as “Could you tell…”, “In what location…”, “Do you know…”, and more. This is vital to resist VL models’ notorious brittleness against linguistic variations Shah et al. (2019); Ray et al. (2019); Kervadec et al. (2021); Whitehead et al. (2020).
Object-Scene Relevance
[…]
4 Experimental Settings
The object hallucination benchmark consists of the validation and test sets of 10 VQA corpora, including NOPE (§3.4) with balanced object-scene relevance proportions. It displays the comparison between incorrectness and object hallucination over various baselines, which serves as a foundation for assessing object hallucination in addition to the standard incorrectness in 10 VL models.
4.1 Datasets
Table 1 describes the data distribution of the dev and test sets of the benchmark. Each set respectively comprises $\sim$30k and $\sim$36k data, maintaining near-balanced proportions of $\mathrm{NegP}$ and $\mathrm{Others}$ data. To ensure the quality of the visual questions in the benchmark, we also analyze the lexical diversity and the fluency of the comprising datasets, which are useful to assert a robust evaluation using questions that are linguistically diverse and coherent. In Figure 6, we show that the datasets whose data construction utilizes automatic question generation, i.e., NOPE and TDIUC, have comparable lexical diversity and fluency to the other datasets, which entirely rely on question generation by human annotators.
For lexical diversity, we employ length-agnostic lexical diversity metrics, i.e., moving average type-token ratio (MATTR) Covington and McFall (2010), measure of textual lexical diversity (MTLD) McCarthy (2005), and hypergeometric distribution diversity (HDD) McCarthy and Jarvis (2007, 2010), and average them. We use LexicalRichness Shen (2021, 2022) v0.5.0333https://pypi.org/project/lexicalrichness/ to calculate these metrics. We also employ a large pre-trained LM GPT-Neo Black et al. (2021) with 2.7B parameters to compute the perplexity of the questions, which is often used as a measure of both lexical diversity Lewis et al. (2017); Tevet and Berant (2021) and fluency Fan et al. (2018); Wang et al. (2019); Cahyawijaya et al. (2021); Anonymous (2023).
Figure 6: Question quality in the benchmark in terms of lexical diversity and fluency.
4.2 Baselines
For the baselines in our benchmark, we employ various vision-language model architectures on the benchmark in both zero-shot & few-shot and fine-tuned fashion. For the fine-tuned setting, we utilize five models: 1) OFA Wang et al. (2022b), which unifies architectures, tasks, and modalities by formulating a unified sequence-to-sequence abstraction via handcrafted instructions to achieve task agnosticism; 2) and 3) BLIP Li et al. (2022), which incorporates two key contributions, i.e., multimodal mixture of encoder-decoder (MED) to operate as either a unimodal encoder, an image-grounded text encoder, or an image-grounded text decoder, and CapFilt as a new dataset bootstrapping method for learning from noisy image-text pairs; 4) ALBEF Li et al. (2021a), which is trained using momentum distillation to improve learning from noisy web data; 5) GIT Wang et al. (2022a), which employs an image encoder and a text decoder pre-trained using a language modeling objective to map the input image to its corresponding description.
For the zero-shot setting, we employ: 1) BLIP-2 Li et al. (2023a), which utilizes a scalable multimodal pre-training method to enable any LLMs to ingest and understand images; 2) and 3) PromptCap Hu et al. (2022), which is trained to generate captions that help downstream LMs answer visual questions; 4) InstructBLIP Dai et al. (2023a), which is an instruction-tuned version of BLIP-2 on various tasks including VQA. We also employ 5) OpenFlamingo Alayrac et al. (2022); Awadalla et al. (2023), which is an open-source version of a large pre-trained VL model specialized in few-shot prompting, in the two-shot setting. Table 2 provides the model and data sizes of the baselines and Appendix H lists the model variants.
Model size # Pre-train images
Zero-shot & Few-shot
[…]
PromptCap 3B 34M
BLIP-2 3.8B 129M
OpenFlamingo 9B $\sim$2.5B
VQA fine-tuned
OFA 929M 34M
BLIP 385M 129M
[…]
ALBEF 628M 14M
[…]
Table 2: VL baseline models in the benchmark.
$\mathbf{Others}$ test (%) $\mathbf{NegP}$ test (%)
Overall Existing datasets NOPE test (§3.4) Overall
Acc. METEOR Acc. METEOR Acc. METEOR $\mathrm{NegP}$ Acc. Acc. METEOR
Zero-shot & few-shot
[…]
PromptCap 32.69 22.66 3.61 2.20 0.42 0.56 1.67 0.99 0.85
BLIP-2 19.84 17.94 4.39 1.49 2.11 1.22 5.25 2.51 1.27
OpenFlamingo 14.29 24.32 0.09 7.96 0.00 0.08 0.02 0.02 1.49
VQA fine-tuned
OFA 29.43 17.06 3.24 4.10 2.75 9.11 8.21 2.84 8.21
BLIP 23.27 12.07 5.95 5.12 1.60 3.63 6.48 2.38 3.90
[…]
ALBEF 16.33 21.87 19.31 26.31 1.86 6.76 8.18 4.98 10.25
[…]
InstructBLIP 40.62 22.55 21.40 13.50 5.08 5.19 17.69 7.99 6.67
Table 3: Weighted model performances on the test set of the benchmark. Errors made on $\mathrm{Others}$ VQA data represent incorrectness, while errors made on $\mathrm{NegP}$ VQA data represent object hallucination. Bold and underline denote the best performances overall and in the group, respectively.
4.3 Evaluation Settings
For both $\mathrm{NegP}$ and $\mathrm{Others}$, we compute accuracy and METEOR Banerjee and Lavie (2005) to measure the performance of vision-language models on the benchmark. While accuracy measures the performance based on an exact match between the generated answer and the ground truth answer, METEOR caters to partial (i.e., unigram) matches by computing a score for this matching using a combination of unigram-precision, unigram-recall, and alignment between the unigrams in the generated answer and ground truth answer. Additionally, for $\mathrm{NegP}$, we employ a rule-based accuracy, referred to as $\mathrm{NegP}$ accuracy, which focuses on determining whether the generated answer is a negative indefinite pronoun (i.e., $\in A^{\mathrm{NegP}}=\{"none","nothing","nowhere","zero","0",\\ "no\;one","nobody","neither"\}$) or not. All scores are computed per task and then the weighted averages according to each task size are retrieved.
5 Results
[…]
Another point that we observe is, GIT outperforms the other baselines on both $\mathrm{NegP}$ and $\mathrm{Others}$ data, as well as manages to surpass much bigger models (e.g., InstructBLIP and Flamingo), showing that GIT is more robust against both object hallucination and general incorrectness, despite being the smallest in size (Table 2) and having a simple architecture. This achievement could be attributed to its substantial number of pre-training images, which is an order of magnitude larger than those of the other baselines. This also aligns with Hoffmann et al. (2022), in which for the same compute budget, a smaller model trained on more data outperforms a larger model trained on fewer data and achieves more optimal performance.
Figure 7: All baselines consistently score lower on $\mathrm{NegP}$ (%$\mathrm{NegP}$ Acc.) than $\mathrm{Others}$ (%Acc.).
6 Analysis and Discussions
6.1 Object hallucination and lexical diversity
Table 3 also show that $\mathrm{NegP}$ performance scores on existing datasets are significantly higher than on NOPE across the metrics, indicating that object hallucination is more likely to occur when the models attempt to solve the questions in NOPE. This is mainly due to the NOPE dataset having a relatively higher lexical diversity compared to the other $\mathrm{NegP}$ corpora, which are mostly composed of VQAv2 and Visual7W (see in Figure 6). This also aligns with the fact that $\mathrm{NegP}$ model performances have a strong negative Pearson correlation with the lexical diversity measures ($r=\{-0.8,-0.66,-0.65,-0.7\}$ for METEOR and HDD, MTLD, MATTR, perplexity) and proves that corpora with higher lexical diversity (e.g., NOPE) provide more challenging $\mathrm{NegP}$ VQA problems to assess object hallucination.
Figure 8: VL models are more prone to object hallucination on lexically diverse $\mathrm{NegP}$ VQA data. Dot size represents dataset size (§4.1).
6.2 Object hallucination and language bias
As shown in Figure 9, among 5 $\mathrm{NegP}$ question types, all VQA-finetuned VL models fail on $\mathrm{NegP}$ questions about color (e.g., “What is the color of…?”), object (e.g., “What is the object beside…?”), and location (e.g., “Where is…?”), while most VL models tend to hallucinate less on $\mathrm{NegP}$ questions about counting (e.g., “How many…?”) and person (e.g., “Who is using…?”). A similar trend is observed for the zero-shot & few-shot baselines. We further inspect these two categories and find out that their answer scopes are of a smaller scope than the others in the training data. For instance, the answers to counting questions are often numbers $\leq$ 5, and the answers to the person questions are often the generic "man", "woman", "person", "people", and others which have fewer variations compared to object types, color names, or absolute and relative places. These facts suggest that existing VL models have a strong language bias KV and Mittal (2020); Niu et al. (2021b); Wu et al. (2022) toward certain question types, which result in acceptable $\mathrm{NegP}$ performances on those question types. Nevertheless, language bias does not solve object hallucination and even might make it worse, due to the VL models having weak visual grounding skills to verify the answer to the visual context, which might lead to errors on both $\mathrm{NegP}$ and $\mathrm{Others}$ questions.
Figure 9: $\mathrm{NegP}$ performance of VQA fine-tuned baselines over different question types.
6.3 Object hallucination and object-scene relevance
As shown in Figure 10, all VQA fine-tuned models perform lower when the object is closely related to the scene compared to when the object is loosely related or unrelated. This indicates that VL models have some degree of understanding $\mathrm{NegP}$ based on the relevance of the object in question with the scene. Although this helps VL models to understand about objects better in some cases, this also causes VL models to hallucinate more on objects that are relevant to the scene Rohrbach et al. (2018); Kayhan et al. (2021); Dai et al. (2023b). On the other hand, the performance on loosely related or unrelated objects tend to be similar, which aligns with the similarity analysis provided in Figure 5. In contrast, for zero-shot & few-shot baselines, the differences between object-scene relevance are less apparent. However, in general, the $\mathrm{NegP}$ scores are also very low, except for BLIP-2, which suggests that most zero-shot models do not have an adequate understanding of $\mathrm{NegP}$.
Figure 10: $\mathrm{NegP}$ performance of (left) zero-shot & few-shot and (right) VQA fine-tuned baselines per object-scene relevance.
7 Conclusion
We have addressed the critical issue of object hallucination in VL models, which has been lacking a general measurement. We have introduced NOPE to assess object hallucination in VL models, investigating the discernment of objects’ non-existence in visual questions by 10 state-of-the-art VL models, alongside their standard performances. Additionally, we have presented a cost-effective and scalable method for generating high-quality synthetic data with over 90% validity to overcome the severe underrepresentation of $\mathrm{NegP}$ cases. Through our comprehensive experiments, we have demonstrated that no VL model is exempt from object hallucination, highlighting their lack of understanding of negative object presence. Furthermore, we have identified lexical diversity, question type, and the relevance of the object to the visual scene as influential factors impacting VL models’ susceptibility to object hallucination. These findings provide valuable insights into the assessment of object hallucination in VL, thereby paving the way for the future development of enhanced VL models.
8 Limitation and Future Work
Evaluation Metrics for Object Hallucination
In this work, we show three metrics to measure object hallucination and incorrectness, i.e., the exact match accuracy, METEOR, and $\mathrm{NegP}$ accuracy. Nevertheless, in some cases, these metrics fail to capture some equivalent answer that has the same semantic meaning. For example, given an $\mathrm{NegP}$ question “Where is the spoon in the picture?” with the corresponding label “Nowhere”, a system that answers with “There is no spoon in the picture” will get 0 scores on these three metrics, despite the answer is actually correct. We argue that the limitation of the existing metrics might hinder further research in alleviating object hallucination and we expect future works to focus on developing better metrics for measuring object hallucination.
Object Hallucination Outside of $\mathbf{NegP}$
Since object hallucination refers to an effect (i.e., generating non-existent objects) and not a cause, our measurement of object hallucination is limited to $\mathrm{NegP}$ cases, in which a VL model unfaithfully infers a supposedly non-existent object as existent in the visual context. For cases where a VL model provides an incorrect answer to $\mathrm{Others}$ VQA, the fine line between misclassification and object hallucination has not yet been defined.
Performances on Full $\mathbf{Others}$ Test Sets
In order to observe the incorrectness of VL models on $\mathrm{Others}$ on various datasets, we compose a balanced set of $\sim$15k data in our dev split and $\sim$18k data in our test split from diverse VQA corpora. Obtaining the full performance on each of the source datasets requires re-running the baselines on the full test sets of each source dataset.
9 Ethics Statement
This research on object hallucination in vision-language models aims to improve the reliability and faithfulness of these models, which have significant applications in various fields such as healthcare and autonomous driving. We acknowledge the potential impact of our findings and commit to promoting responsible and ethical use of these models. We recognize that such models have the potential to perpetuate biases and stereotypes, and we have taken steps to mitigate this risk. For instance, we ensured that the synthetic data used in this study was generated in a manner that respects privacy and does not perpetuate biases or stereotypes. Furthermore, we recognize the importance of transparency and accountability in the development and use of these models. Therefore, we commit to sharing our findings and methodologies openly and making them accessible to the wider research community. We also acknowledge that these models can have unintended consequences and commit to ongoing monitoring and evaluation of their impact. Finally, we recognize that the development and use of these models must be guided by ethical principles that prioritize human well-being and social responsibility. We are committed to upholding these principles and contributing to the development of responsible and ethical practices in the field of vision-language modeling.
Appendix A Prompt Templates
We provide the prompt templates and examples for the generate-from-scratch and list-then-rewrite methods in Table 4 and Table 5, respectively.
Template 1
Create a question beginning with "<INTERROGATIVE_WORD>" from this image caption: "<IMAGE_CAPTION>" with an answer of "<ANSWER>".
Example: Create a question beginning with "who" from this image caption: "This image consists of an airplane in the air. On which, we can see the text. In the background, there is sky." with an answer of "nobody".
Generated question: Who is in the airplane in this image?
Template 2
Turn the text "<IMAGE_CAPTION>" into a "<INTERROGATIVE_WORD>" question with an answer "<ANSWER>":
Example: Turn the text "This is an outside view. Here I can see the cars on the ground and there are two persons standing. In the background there is a tent, shelter and trees. At the top of the image there is the sky. At the bottom there is grass." into a "how many" question with an answer "none":
Generated question: How many buildings are visible in the outside view?
Template 3
Given an answer "<ANSWER>", generate an appropriate question starting with "<INTERROGATIVE_WORD>" from the text "<IMAGE_CAPTION>".
Example: Given an answer "nowhere", generate an appropriate question starting with "where" from the text "In this picture, we see a woman. She is staring at something and she might be posing for the photo. We see a nail is pierced on her face. She might be taking bath. In the background, we see a wall and a white color object which looks like a bathtub.".
Generated question: Where is the ideal location to put a nail on a person’s face while they are taking a bath?
Template 4
Write a question that begins with "<INTERROGATIVE_WORD>" and is answered by "<ANSWER>" based on the text "<IMAGE_CAPTION>".
Example: Write a question that begins with "what" and is answered by "nothing" based on the text "In this image I can see food item in a cooking vessel, which is on the stove. In the background there are some objects.".
Generated question: What can you see in the background of the image?
Template 5
Rewrite the statement "<IMAGE_CAPTION>" as a question that starts with interrogative phrase "<INTERROGATIVE_WORD>" and has an answer of "<ANSWER>".
Example: Rewrite the statement "In the image I can see food item in the bowl. The background of the image is blurred." as a question that starts with interrogative phrase "which" and has an answer of "neither".
Generated question: Which part of the image is not blurred and which object is visible in the bowl?
Table 4: Prompt templates used in the generate-from-scratch3.1) and their corresponding examples.
Object Listing Template
Description:
"<IMAGE_CAPTION>"
List 10 tangible objects that are closely related to the description, but are not mentioned there. Also avoid these objects: <IMAGE_LABELS>.
Example:
Description:
"In this picture we can see some food products in the glass jars."
List 10 tangible objects that are closely related to the description, but are not mentioned there. Also avoid these objects: Vegetable.
Response:
1. Spices, 2. Canned fruits, 3. Nuts, 4. Sugar, 5. Coffee beans, 6. Honey, 7. Pasta, 8. Chocolates, 9. Tea leaves, 10. Jams and preserves
Question Rewriting Template
Question templates:
<REFERENCE_QUESTIONS_BASED_ON_FIRST_TURN_ANSWER>
Paraphrase each of these <NUMBER> questions while keeping the sentence as a question with a question mark.
Example:
Question templates:
1. Who is using the object above the spices?
2. What object is located under the canned fruits?
3. Who is holding the object on the right of the nuts?
4. What is the color of the sugar?
5. Who is using the coffee beans?
6. What object is located on the left of the honey?
7. What object is located on the left of the pasta?
8. What object is located on the left of the chocolates?
9. What is the color of the tea leaves?
10. Where is the jams and preserves located in the image?
Paraphrase each of these 10 questions while keeping the sentence as a question with a question mark.
Generated question:
1. Who is the user of the spices in the picture?
2. Which product can be found below the canned fruits?
3. Who is the individual holding the object to the right of the nuts?
4. Can you tell me the color of the sugar in the glass jars?
5. Who might be utilizing the coffee beans shown in the image?
6. What is the item on the right side of the honey?
7. What comes up on the left of the pasta in the photograph?
8. What is situated on the left of the chocolates in the picture?
9. What is the color of the tea leaves displayed in the glass jars?
10. Where exactly can the jams and preserves be found in this image?
Table 5: Prompt templates used in the list-then-rewrite3.1) and their corresponding examples.
Appendix B Reference Question Templates
Table 6 presents the pool of question templates used to automatically build the reference questions for the list-then-rewrite in §3.1.
No Question template
$\mathbf{NegP}$ answer
1
What is the color of the <OBJECT>?
none / nothing
2
What object is located above / under / on the left of / on the right of the <OBJECT>?
none / nothing
3
Where is the <OBJECT> located in the image?
nowhere
4
How many <OBJECT> are there in the image?
zero / 0 / none
5
Who is holding / using the <OBJECT>?
no one / nobody
6
Who is holding / using the object above / under / on the left of / on the right of the <OBJECT>?
no one / nobody
Table 6: Question templates utilized to construct the reference questions for the question rewriting step in the list-then-rewrite prompting methodology in §3.1.
Appendix C Automatic Validation Methodologies of $\mathbf{NegP}$ VQA Data Generation
Generate-from-scratch
To ensure the validity of $q_{i}$, we use a model fine-tuned on natural language inference (NLI) to determine whether a generated question $q_{i}$ and answer $a_{i}$ pair (i.e., hypothesis) logically entails its corresponding image caption $c_{i}$ (i.e., premise). We also utilize a fine-tuned binary classifier to determine whether a generated question $q_{i}$ and answer $a_{i}$ pair fits a given visual context $v_{i}$. If the question $q_{i}$ and answer $a_{i}$ pair is true (entailment) or undetermined (neutral) given $c_{i}$ as well as matches with $v_{i}$, then the generated question $q_{i}$ is judged as valid by the automatic validation.
List-then-rewrite
For the automatic validation of a listed object $o_{i,j}$, we extract lemmatized noun tokens from its corresponding image caption $c_{i}$ and obtain the object names from $l_{i}$ as the objects present in $v_{i}$. If $o_{i,j}$ does not match with any of the extracted objects, then $o_{i,j}$ is a valid non-existent object. For the automatic validation of a generated question $q_{i,j}$, if $q_{i,j}$ does not contradict its respective reference question $r_{i,j}$, then the generated question $q_{i,j}$ is considered valid.
Appendix D Implementation Details of $\mathbf{NegP}$ VQA Data Generation
We implement §3.1 with the following LLMs that employ: 1) multi-task prompted fine-tuning, i.e., BLOOMZ Muennighoff et al. (2022) and T0 Sanh et al. (2022); 2) instruction meta-learning, i.e., OPT-IML Iyer et al. (2022); 3) synthetic self-instruct, i.e., Alpaca Wang et al. (2022c); 4) instruction Wei et al. (2022a) and chain-of-thought fine-tuning Wei et al. (2022b), i.e., FLAN T5 and FLAN Alpaca Chung et al. (2022); 5) multi-task instruction pre-training, i.e., ChatGLM Zeng et al. (2023); 6) conversation-style instruction tuning and reinforcement learning with human feedback (RLHF) Christiano et al. (2017); Stiennon et al. (2020), i.e., ChatGPT (GPT-3.5). More details are presented in Table 7.
No Model Size References Access
1 BLOOMZ (3B) 3B Muennighoff et al. (2022); Scao et al. (2022) https://huggingface.co/bigscience/bloomz-3b
2 BLOOMZ (7.1B) 7.1B Muennighoff et al. (2022); Scao et al. (2022) https://huggingface.co/bigscience/bloomz-7b1
3 T0 3B Sanh et al. (2022) https://huggingface.co/bigscience/T0_3B
4 OPT-IML 1.3B Iyer et al. (2022); Zhang et al. (2022) https://huggingface.co/facebook/opt-iml-max-1.3b
5 Alpaca 7B Wang et al. (2022c); Touvron et al. (2023) https://huggingface.co/chavinlo/alpaca-native
6 FLAN T5 XL 3B Chung et al. (2022); Raffel et al. (2020) https://huggingface.co/google/flan-t5-xl
7 FLAN T5 XXL 11B Chung et al. (2022); Raffel et al. (2020) https://huggingface.co/google/flan-t5-xxl
8 FLAN Alpaca XL 3B Chung et al. (2022); Wang et al. (2022c) https://huggingface.co/declare-lab/flan-alpaca-xl
9 ChatGLM 6B Zeng et al. (2023); Du et al. (2022) https://huggingface.co/THUDM/chatglm-6b
10 ChatGPT 175B - https://platform.openai.com/docs/models/gpt-3-5
Table 7: Instruction-tuned LLMs used in Appendix D.
We utilize Open Images v7 as our image captioning dataset $\mathcal{D}_{cap}$ with respect to the provided splits. For automatic validation with NLI, we use the RoBERTa model fine-tuned on various NLI corpora that achieves the best performance on the Adversarial NLI benchmark Nie et al. (2020). [^4] https://huggingface.co/ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli For automatic validation with image-QA pair classification, we build a simple CLIP-based Radford et al. (2021) binary classifier. We provide the details in Appendix D.1. For the list-then-rewrite method, we use $m=10$.
D.1 Image-QA Pair Classification
To construct a model for our image-QA pair classification, we construct a balanced image-QA corpus using $\mathrm{NegP}$ and $\mathrm{Others}$ VQA data randomly selected from 9 existing VQA datasets, i.e., VQAv2 (Balanced Real) Antol et al. (2015), AdVQA Sheng et al. (2021), VizWiz Gurari et al. (2018, 2019), TextVQA Singh et al. (2019), R-VQA Lu et al. (2018), Visual7W Zhu et al. (2016), TDIUC Kafle and Kanan (2017), VQA-Rephrasings Shah et al. (2019), and VQAv1 (Abstract Scenes) Antol et al. (2015).
For the image-QA pairs from the $\mathrm{NegP}$ VQA data, we assign a binary label of 1 (valid), which means that the QAs correctly fit the corresponding images as valid pairs. For the $\mathrm{Others}$ VQA data, we replace the $\mathrm{Others}$ ground truth answers with $\mathrm{NegP}$ answers $\in A^{\mathrm{NegP}}$ to make the invalid image-QA pairs (a binary label of 0). We split the corpus into 6k training, 2k validation, and 2k test set.
Using this corpus, we train a simple classifier with one hidden layer on top of a frozen CLIP Radford et al. (2021). We leverage the image-text alignment learned by CLIP Radford et al. (2021), which has been pre-trained on 400M image-text pairs using contrastive learning, to extract the image features of the images and the textual features of their question-answer counterparts. We simply concatenate both image and text features, then input them into the classifier. Our image-QA pair classifier yields an F1-score of 91.29% on the test set.
Appendix E Human Evaluation Category Examples
We provide the human evaluation categories (§3.2) in Figure 11.
Figure 11: Examples of the human evaluation judgments for the generate-from-scratch prompting method in §3.2.
Appendix F Automatic Validation Results of $\mathrm{NegP}$ VQA Data Generation
Figure 12: Automatic validation results on 1000 $\mathrm{NegP}$ questions generated using generate-from-scratch3.1) over five prompt templates.
Figure 13: Human evaluation results on $\mathrm{NegP}$ questions generated by ChatGPT using generate-from-scratch3.1). The Y-axis denotes the verdict from the automatic validators, i.e., caption-QA and image-QA entailment models.
Generate-from-scratch
Figure 12 shows the proportions of valid generated $\mathrm{NegP}$ VQA data using 10 instruction-tuned LLMs listed in Appendix D over five different prompt templates, where each model generates 1k questions per template. The prompt templates are provided in Appendix A. The result shows that only $\sim$25of the generated questions by the best-performing model, ChatGPT, are valid according to the automatic validation, while other models’ valid generated questions range from 6%-23%. This indicates that the task of $\mathrm{NegP}$ question generation is more complex and difficult than the instructions used to fine-tune the LLMs.
Next, we conduct a human evaluation on randomly selected 240 generated questions (i.e., 60 for each category in §3.2) by ChatGPT, which is the best-performing model. We ask 3 human experts to judge each generated question and answer pair into one of the five options defined in §3.2. Figure 13 demonstrates the result of our human evaluation. The result shows that automatic validation judgments do not agree with the human judgments on a considerable amount of the data, even for simple valid/invalid classification, the automatic validation judgments misclassify 27%-50% of the subsets. From this result, we can conjecture that our automatic validation approach is not effective at verifying whether the generated $\mathrm{NegP}$ questions are valid or invalid and that the generate-from-scratch prompting method is not reliable and fails to elicit the LLMs’ understanding of the task.
List-then-rewrite
Instruction-tuned LLM % Valid objects % Valid objects & questions
FLAN T5 XL
11
10
FLAN T5 XXL
5
17
Alpaca
44
53
FLAN Alpaca XL
25
11
ChatGLM
84
44
ChatGPT
99
98
Table 8: Automatic validation results on 100 $\mathrm{NegP}$ questions generated using list-then-rewrite3.1).
Figure 14: Human evaluation results on $\mathrm{NegP}$ questions generated by ChatGPT using list-then-rewrite3.1).
The automatic validation results on 100 generated questions (i.e., with the category proportion of 50, 35, and 15, respectively) by list-then-rewrite are provided in Table 8. The best-performing model, ChatGPT, yields 98% valid questions with a valid non-existent object according to the automatic validation judgments, which is a huge improvement compared to generate-from-scratch. Similarly, Alpaca and ChatGLM also experience the same increase in validity (albeit not as significant), while the FLAN family models deteriorate due to their inability to handle lists inside the instructions, thus forcing them to respond with only one object instead of 10 objects (§D).
Our human evaluation on 300 generated questions by ChatGPT (presented in Figure 14) also proves that, when we omit the question generation on the wrong object, we can achieve around 90% high-quality $\mathrm{NegP}$ questions generated by the list-the-rewrite method. However, this method would benefit from the establishment of a more suitable penalizing method to filter out the generated questions that are inconsistent with the image captions.
Appendix G Question Diversity of Existing VQA Datasets
We provide the illustrations of question diversity of existing VQA datasets: VQAv2 dataset Antol et al. (2015) which utilizes a manual data generation method (presented in Figure 14(a)) and VQA-Rephrasings dataset Shah et al. (2019) which utilizes an automatic data generation method (presented in Figure 14(b)).
(a) VQAv2 dataset Antol et al. (2015)
[…]
Figure 15: Question diversity of existing datasets. The figures are taken from the respective original papers.
Appendix H Baselines in NOPE Benchmark
The variant details of the baselines used in NOPE are presented in Table 9.
No Model References Access
Zero-shot & Few-shot
[…]
Hu et al. (2022)
https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-base
2 PromptCap
Hu et al. (2022)
https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-3b
3 BLIP-2
Li et al. (2023a)
https://huggingface.co/Salesforce/blip2-opt-2.7b
4 OpenFlamingo
Alayrac et al. (2022); Awadalla et al. (2023)
https://huggingface.co/OpenFlamingo/OpenFlamingo-9B
5 InstructBLIP
Dai et al. (2023a)
https://huggingface.co/Salesforce/instructblip-flan-t5-xl
VQA fine-tuned
1 OFA
Wang et al. (2022b)
https://huggingface.co/OFA-Sys/ofa-huge-vqa
2 BLIP
Li et al. (2022)
https://huggingface.co/Salesforce/blip-vqa-base
[…]
Li et al. (2022)
https://huggingface.co/Salesforce/blip-vqa-capfilt-large
4 ALBEF
Li et al. (2021a)
https://github.com/salesforce/ALBEF#download#Finetuned-checkpoint-for-VQA
[…]
Wang et al. (2022a)
https://huggingface.co/microsoft/git-large-vqav2
Table 9: Variant details of the baselines in NOPE (§4.2).
Appendix I Examples of Object Hallucination in NOPE
We list the examples of object hallucination from the dev set of NOPE in Table 10.
ID Object-scene relevance Visual context Question-answer
390 Related [Uncaptioned image] Question
can you point out the location of the dog collar in the image?
GT answer
nowhere
PromptCap yes
OpenFlamingo The dog collar is located on the dog’s neck.
BLIP-2 yes, it is on the dog’s collar
OFA
no
BLIP
no
ALBEF dog’s neck
[…]
no
InstructBLIP
no
822 Related [Uncaptioned image] Question
how many menu cards have been captured in the image?
GT answer
none
PromptCap 1
OpenFlamingo 10
BLIP-2
none
OFA 8
BLIP six
ALBEF 2,3
[…]
0
InstructBLIP
0
982 Related [Uncaptioned image] Question
where can the clouds be spotted in the image?
GT answer
nowhere
PromptCap in the sky
OpenFlamingo On the head of a goose.
BLIP-2 the clouds are in the sky
OFA the sky
BLIP yes
ALBEF in the snow.
[…]
no
InstructBLIP sky
9165 Partially related [Uncaptioned image] Question
who can you see using the fishing rod?
GT answer
nobody
PromptCap a gray van
OpenFlamingo The owner of this Dodge B250 van.
BLIP-2 the guy in the back of the van
OFA
no 1
BLIP
no idea
ALBEF man dancing
[…]
no
InstructBLIP
no one
10135 Unrelated [Uncaptioned image] Question
which color is the pillow in the image?
GT answer
nothing
PromptCap blue
OpenFlamingo blue
BLIP-2 blue
OFA black
BLIP red and white
ALBEF red black white
[…]
InstructBLIP white
Table 10: Examples of object hallucination in the dev set of NOPE. The hallucinated answers are shown in pink.

### v2

@@
 Zero-shot & Few-shot
-[…]
+PromptCapBASE 696M 34M
 PromptCap 3B 34M
@@
 BLIP 385M 129M
-[…]
+BLIPCapFilt-L 385M 129M
 ALBEF 628M 14M
-[…]
+GITLARGE 347M 1.4B
 […]
@@
 Zero-shot & few-shot
-[…]
+PromptCapBASE 30.18 21.45 2.87 3.05 0.21 0.29 0.95 0.68 0.78
 PromptCap 32.69 22.66 3.61 2.20 0.42 0.56 1.67 0.99 0.85
@@
 BLIP 23.27 12.07 5.95 5.12 1.60 3.63 6.48 2.38 3.90
-[…]
+BLIPCapFilt-L 23.28 12.08 5.95 5.12 1.60 3.61 6.47 2.37 3.88
 ALBEF 16.33 21.87 19.31 26.31 1.86 6.76 8.18 4.98 10.25
-[…]
+GITLARGE 41.00 21.75 34.89 20.43 4.00 5.90 17.92 9.51 8.49
 InstructBLIP 40.62 22.55 21.40 13.50 5.08 5.19 17.69 7.99 6.67
@@
 5 Results
-[…]
+We present the results on the test set of the benchmark in Table 3.
+Examples of object hallucination are in Appendix I.
+While the VQA-finetuned baselines are slightly better at $\mathrm{NegP}$ and comparable to the zero-shot & few-shot baselines on $\mathrm{Others}$, as in Figure 7, we observe that all zero-shot and VQA-finetuned baselines notably perform much worse on $\mathrm{NegP}$ tasks that $\mathrm{Others}$ with the averaged discrepancies of $\pm$22and $\pm$18accuracy, respectively.
+This demonstrates that all baselines are more vulnerable and susceptible to object hallucination than standard incorrectness.
+In addition, less incorrectness does not entail less object hallucination.
+For instance, PromptCapBASE, PromptCap, and BLIP have lower scores on $\mathrm{NegP}$ than ALBEF despite outperforming it on $\mathrm{Others}$ setting.
+It also means that existing evaluations that solely utilize $\mathrm{Others}$ cases cannot effectively capture the models’ risk of object hallucination.
 Another point that we observe is, GIT outperforms the other baselines on both $\mathrm{NegP}$ and $\mathrm{Others}$ data, as well as manages to surpass much bigger models (e.g., InstructBLIP and Flamingo), showing that GIT is more robust against both object hallucination and general incorrectness, despite being the smallest in size (Table 2) and having a simple architecture.
@@
 We are committed to upholding these principles and contributing to the development of responsible and ethical practices in the field of vision-language modeling.
+Acknowledgements
+This work has been partially funded by PhD Fellowship Award, the Hong Kong University of Science and Technology; PF20-43679 Hong Kong PhD Fellowship Scheme, Research Grant Council, Hong Kong; and the National Research Foundation, Singapore under its AI Singapore Programme.
 Appendix A Prompt Templates
@@
 Table 6 presents the pool of question templates used to automatically build the reference questions for the list-then-rewrite in §3.1.
-No Question template
-$\mathbf{NegP}$ answer
-1
-What is the color of the <OBJECT>?
-none / nothing
-2
-What object is located above / under / on the left of / on the right of the <OBJECT>?
-none / nothing
-3
-Where is the <OBJECT> located in the image?
-nowhere
-4
-How many <OBJECT> are there in the image?
-zero / 0 / none
-5
-Who is holding / using the <OBJECT>?
-no one / nobody
-6
-Who is holding / using the object above / under / on the left of / on the right of the <OBJECT>?
-no one / nobody
+No Question template $\mathbf{NegP}$ answer
+1 What is the color of the <OBJECT>? none / nothing
+2 What object is located above / under / on the left of / on the right of the <OBJECT>? none / nothing
+3 Where is the <OBJECT> located in the image? nowhere
+4 How many <OBJECT> are there in the image? zero / 0 / none
+5 Who is holding / using the <OBJECT>? no one / nobody
+6 Who is holding / using the object above / under / on the left of / on the right of the <OBJECT>? no one / nobody
 Table 6: Question templates utilized to construct the reference questions for the question rewriting step in the list-then-rewrite prompting methodology in §3.1.
@@
 Instruction-tuned LLM % Valid objects % Valid objects & questions
-FLAN T5 XL
-11
-10
-FLAN T5 XXL
-5
-17
-Alpaca
-44
-53
-FLAN Alpaca XL
-25
-11
-ChatGLM
-84
-44
-ChatGPT
-99
-98
+FLAN T5 XL 11 10
+FLAN T5 XXL 5 17
+Alpaca 44 53
+FLAN Alpaca XL 25 11
+ChatGLM 84 44
+ChatGPT 99 98
 Table 8: Automatic validation results on 100 $\mathrm{NegP}$ questions generated using list-then-rewrite3.1).
@@
 (a) VQAv2 dataset Antol et al. (2015)
-[…]
+(b) VQ2A dataset Changpinyo et al. (2022)
 Figure 15: Question diversity of existing datasets. The figures are taken from the respective original papers.
@@
 The variant details of the baselines used in NOPE are presented in Table 9.
-No Model References Access
-Zero-shot & Few-shot
-[…]
-Hu et al. (2022)
-https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-base
-2 PromptCap
-Hu et al. (2022)
-https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-3b
-3 BLIP-2
-Li et al. (2023a)
-https://huggingface.co/Salesforce/blip2-opt-2.7b
-4 OpenFlamingo
-Alayrac et al. (2022); Awadalla et al. (2023)
-https://huggingface.co/OpenFlamingo/OpenFlamingo-9B
-5 InstructBLIP
-Dai et al. (2023a)
-https://huggingface.co/Salesforce/instructblip-flan-t5-xl
-VQA fine-tuned
-1 OFA
-Wang et al. (2022b)
-https://huggingface.co/OFA-Sys/ofa-huge-vqa
-2 BLIP
-Li et al. (2022)
-https://huggingface.co/Salesforce/blip-vqa-base
-[…]
-Li et al. (2022)
-https://huggingface.co/Salesforce/blip-vqa-capfilt-large
-4 ALBEF
-Li et al. (2021a)
-https://github.com/salesforce/ALBEF#download#Finetuned-checkpoint-for-VQA
-[…]
-Wang et al. (2022a)
-https://huggingface.co/microsoft/git-large-vqav2
 Table 9: Variant details of the baselines in NOPE (§4.2).
@@
 ID Object-scene relevance Visual context Question-answer
-390 Related [Uncaptioned image] Question
-can you point out the location of the dog collar in the image?
-GT answer
-nowhere
+390 Related [Uncaptioned image] Question can you point out the location of the dog collar in the image?
+GT answer nowhere
 PromptCap yes
@@
 BLIP-2 yes, it is on the dog’s collar
-OFA
-no
-BLIP
-no
+OFA no
+BLIP no
 ALBEF dog’s neck
-[…]
-no
-InstructBLIP
-no
-822 Related [Uncaptioned image] Question
-how many menu cards have been captured in the image?
-GT answer
-none
+GITLARGE no
+InstructBLIP no
+822 Related [Uncaptioned image] Question how many menu cards have been captured in the image?
+GT answer none
 PromptCap 1
 OpenFlamingo 10
-BLIP-2
-none
+BLIP-2 none
 OFA 8
@@
 ALBEF 2,3
-[…]
-0
-InstructBLIP
-0
-982 Related [Uncaptioned image] Question
-where can the clouds be spotted in the image?
-GT answer
-nowhere
+GITLARGE 0
+InstructBLIP 0
+982 Related [Uncaptioned image] Question where can the clouds be spotted in the image?
+GT answer nowhere
 PromptCap in the sky
@@
 ALBEF in the snow.
-[…]
-no
+GITLARGE no
 InstructBLIP sky
-9165 Partially related [Uncaptioned image] Question
-who can you see using the fishing rod?
-GT answer
-nobody
+9165 Partially related [Uncaptioned image] Question who can you see using the fishing rod?
+GT answer nobody
 PromptCap a gray van
@@
 BLIP-2 the guy in the back of the van
-OFA
-no 1
-BLIP
-no idea
+OFA no 1
+BLIP no idea
 ALBEF man dancing
-[…]
-no
-InstructBLIP
-no one
-10135 Unrelated [Uncaptioned image] Question
-which color is the pillow in the image?
-GT answer
-nothing
+GITLARGE no
+InstructBLIP no one
+10135 Unrelated [Uncaptioned image] Question which color is the pillow in the image?
+GT answer nothing
 PromptCap blue
@@
 ALBEF red black white
-[…]
+GITLARGE blue
 InstructBLIP white